The base of a triangular sail is 14 m shorter than its height. The area is 120 m^2. Find the base and height of the sail.
I'm not sure what the formula would be for this.

Answer by imapoet_2003(3)

The formula for the area of a triangle is a=1/2b*h. We know that the base is 14m shorter than the height. This is represented by b=h-14. Knowing this, the problem can be set up as follows: a=1/2(h-14)*h. Since we know that the area of this triangle is 120m^2, the formula you are looking at is 120=1/2(h-14)*h. Hope this helps.